OpenStudy (anonymous):

If 2^x=10, then 2^2x= a)20 b)40 c)80 d)100 e)200

OpenStudy (zzr0ck3r):

hint: \(2^{2x}=(2^x)^2\)

OpenStudy (anonymous):

@zzr0ck3r, is it 100? since 2^x= 10 and (2^x)^2=2^2x, (2^x)^2=(10)^2=100?

OpenStudy (anonymous):

100 is correct

OpenStudy (anonymous):

okay thanks

OpenStudy (zzr0ck3r):

yep. :)